OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 6(x - 2) alt= 6x - 12

OpenStudy (cwrw238):

this is the distributive property the 6 multiplies each term in the bracket: 6(x - 2) = 6*x - 6*2 = 6x - 12